A sample of K2CO3 weighing 27.6g was treated by a series of reagents so as to convert all of its carbon to potassium zinc ferrocyanide [also known as dipotassium trizinc hexacyanoferrate(II)] K2Zn3[Fe(CN)6]2. How many grams of the product were obtained?

1
Expert's answer
2022-01-10T13:58:05-0500

Molar mass of K2CO3= 138.025

27.6/138.025= 0.2

Molar mass of K2Zn3[Fe(CN)6]2= 698.3224

= 0.2×698.3224=139.7g